x+7y=13 Find the slope and the y-int. I can't find it.
Answer by Boreal(15235)

write in form y=mx+b
that would be 7y=-x+13 and dividing by 7 y=-(1/7)x+13/7
the slope is the number in front of the x or -1/7
the y-intercept is the number at the end, 13/7